# Generative Modeling through Spectral Analysis of Koopman Operator

### v1

@@ 4.1 Data-Driven Error Bound Analysis
Let $\rho_{t}=\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}$ and $f_{t}=\rho_{t}-1$. Recall the KL divergence $\mathrm{KL}(\mu_{t}\|\pi)=\int\rho_{t}\log\rho_{t}\,\mathrm{d}\pi$. By differentiation under the invariant measure $\pi$, see [chewi2023log, chewi2020svgd] for more details), we obtain
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\big\langle\nabla f_{t},\,\nabla\mathcal{K}_{r}f_{t}\big\rangle_{\pi}=-\langle f_{t},\mathcal{L}\mathcal{K}_{r}f_{t}\rangle_{\pi}=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi},$ (4.3)
where we used the Dirichlet form identity $\langle\nabla f,\nabla g\rangle_{\pi}=\langle f,\mathcal{L}g\rangle_{\pi}$ and the operator relation (4.1).
Expanding the right-hand side of (4.3), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}=-\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2},$ (4.4)
@@
Plugging into (4.4), we have
@@
where we used $\chi^{2}(\mu_{t}\|\pi)=\|f_{t}\|_{L^{2}_{\pi}}^{2}$. Applying the inequality $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$ to the above inequality, we obtain
@@
which is exactly (4.2). ∎
Remark 4.4.
@@
Remark 4.5.
Proposition 4.3 reveals that the idealized KSWGD (using exact but truncated eigenpairs) inherits the scale-free exponential convergence rate $e^{-t}$ from LAWGD, independent of the Poincaré constant. However, the flow converges to an approximated stationary state with residual error bounded by $\eta_{r}^{2}$. This truncation errors quantifies the steady-state deviation caused by neglecting spectral components beyond the first $r$ modes, and vanishes as $r\to\infty$ by the completeness of the eigenfunction basis, which recovers exact LAWGD.
In practice, the eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{r}$ can be estimated from data using EDMD or kernel-EDMD, which gives approximated eigenpairs $\{(\hat{\lambda}_{i},\hat{\phi}_{i})\}_{i=1}^{r}$ and constructs truncated inverse operator $\widehat{\mathcal{K}}_{r}=\sum_{i=1}^{r}\frac{1}{\hat{\lambda}_{i}}\langle\cdot,\hat{\phi}_{i}\rangle_{\pi}\,\hat{\phi}_{i}$ as in (3.2). Therefore, instead of the exact relation $\mathcal{L}\mathcal{K}_{r}=\Pi_{r}$, the data-driven operator satisfies a perturbed relation
$\mathcal{L}\widehat{\mathcal{K}}_{r}f=\Pi_{r}f+\delta_{r}(f),\quad\forall f\in L^{2}_{\pi},$ (4.5)
Assumption 4.6 (Controlled Operator Approximation Error).
There exists a constant $0<\varepsilon_{r}<<1$ such that for all $t\geq 0$,
Remark 4.7.
Theorem 4.8 (Error Bound Analysis).
Then
$\mathrm{KL}(\mu_{t}\|\pi)\;\leq\;e^{-(1-\varepsilon_{r})t}\,\mathrm{KL}(\mu_{0}\|\pi)\;+\;\frac{\eta_{r}^{2}}{1-\varepsilon_{r}}\Big(1-e^{-(1-\varepsilon_{r})t}\Big).$ (4.6)
@@
Following the same derivation as Proposition 4.3 up to the dissipation identity in (4.3), but now using the perturbed relation $\mathcal{L}\widehat{\mathcal{K}}_{r}=\Pi_{r}+\delta_{r}$ as in (4.5) instead:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\langle f_{t},\mathcal{L}\widehat{\mathcal{K}}_{r}f_{t}\rangle_{\pi}=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}-\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}.$ (4.7)
@@
$\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}\leq|\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}|\leq\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$ (4.8)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}+\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$ (4.9)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-(1-\varepsilon_{r})\|f_{t}\|_{L^{2}_{\pi}}^{2}+\eta_{r}^{2}=-(1-\varepsilon_{r})\chi^{2}(\mu_{t}\|\pi)+\eta_{r}^{2}.$ (4.10)
@@
Remark 4.9.
4.2 Discrete-Time Analysis via Approximate Gradient Flow
While our analysis in Proposition 4.3 and Theorem 4.8 establishes exponential convergence for the continuous-time dynamics, practical implementation relies on the discrete-time update rule (Algorithm 1). To rigorously bridge this gap, we adopt the Approximate Gradient Flow (AGF) framework (See [fujisawa2025on] for more details).
$\alpha=1-\varepsilon_{r}\quad\text{and}\quad\beta=\eta_{r}^{2},$
where $\alpha$ represents the dissipation rate of the update direction and $\beta$ represents the bias floor. By adapting the discrete-time analysis of [fujisawa2025on] to our $\chi^{2}$-geometry, we can directly translate our continuous-time guarantees to the discrete setting.
Corollary 4.10 (Discrete-Time Linear Convergence).
Consider the discrete KSWGD update $\mu_{t+1}=(\mathrm{I}-h\widehat{v}_{t})\#\mu_{t}$ with step size $h>0$ where $\widehat{v}_{t}=-\nabla\widehat{\mathcal{K}}_{r}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)$ [pmlr-v162-salim22a]. Under the conditions of Theorem 4.8 and assuming $\mu_{t}$ is sufficiently close to $\pi$ (i.e., $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$), the discrete iterations satisfy:
$\mathrm{KL}(\mu_{t+1}\|\pi)\leq(1-\alpha h)\mathrm{KL}(\mu_{t}\|\pi)+h\beta+\mathcal{O}(h^{2}).$ (4.11)
@@ 4.2 Discrete-Time Analysis via Approximate Gradient Flow
For any velocity field $v_{t}$, the one-step change in KL satisfies [ambrosio2008gradient, jordan1998variational]
@@
From the continuous-time dissipation inequality (4.10) (or its discrete analogue established in Theorem 4.8), the inner product term is bounded by
@@
Since $\chi^{2}(\mu_{t}\|\pi)\geq 2\mathrm{KL}(\mu_{t}\|\pi)$ holds locally whenever $\mu_{t}$ is sufficiently close to $\pi$, we obtain the desired linear recursion

### v3

@@ Appendix B Theory
Let $\rho_{t}=\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}$ and $f_{t}=\rho_{t}-1$. Recall the KL divergence $\mathrm{KL}(\mu_{t}\|\pi)=\int\rho_{t}\log\rho_{t}\,\mathrm{d}\pi$. By differentiation under the invariant measure $\pi$, see (Chewi, 2023; Chewi et al., 2020) for more details), we obtain
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\big\langle\nabla f_{t},\,\nabla\mathcal{K}_{r}f_{t}\big\rangle_{\pi}=-\langle f_{t},\mathcal{L}\mathcal{K}_{r}f_{t}\rangle_{\pi}=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi},$ (18)
where we used the Dirichlet form identity $\langle\nabla f,\nabla g\rangle_{\pi}=\langle f,\mathcal{L}g\rangle_{\pi}$ and the operator relation (10).
Expanding the right-hand side of (18), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}=-\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2},$ (19)
@@
Plugging into (19), we have
@@
where we used $\chi^{2}(\mu_{t}\|\pi)=\|f_{t}\|_{L^{2}_{\pi}}^{2}$.
Applying the inequality $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$ to the above inequality, we obtain
@@
which is exactly (11). ∎
Remark B.1.
@@
B.2 Proof of Theorem 4.8
@@
Following the same derivation as Proposition 4.3 up to the dissipation identity in (18), but now using the perturbed relation $\mathcal{L}\widehat{\mathcal{K}}_{r}=\Pi_{r}+\delta_{r}$ as in (12) instead:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\langle f_{t},\mathcal{L}\widehat{\mathcal{K}}_{r}f_{t}\rangle_{\pi}=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}-\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}.$ (20)
@@
$\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}\leq|\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}|\leq\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$ (21)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}+\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$ (22)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-(1-\varepsilon_{r})\|f_{t}\|_{L^{2}_{\pi}}^{2}+\eta_{r}^{2}=-(1-\varepsilon_{r})\chi^{2}(\mu_{t}\|\pi)+\eta_{r}^{2}.$ (23)
@@
B.3 Proof of Corollary 4.10
@@
For any velocity field $v_{t}$, the one-step change in KL satisfies (Ambrosio et al., 2008; Jordan et al., 1998)
@@
From the continuous-time dissipation inequality (23) (or its discrete analogue established in Theorem 4.8), the inner product term is bounded by
@@
Since $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$ holds locally whenever $\mu_{t}$ is sufficiently close to $\pi$, we obtain the desired linear recursion